# SiriusBI: Building End-to-End Business Intelligence Enhanced by Large Language Models

### v1

@@ 5.3.1. Evaluation of MRD-Q
In this part, we present the ablation study on dialogue analysis and intention querying using the MRD dataset/ The experimental results are displayed in Table 4.
Based on the experimental results, we can observe that, both the SiriusBI one-step and SiriusBI two-step heavily rely on the dialogue analysis and the intention querying modules. Specifically, when both modules are removed (without MRD-Q), the performance of SiriusBI declines significantly, with drops of 34.37% and 35.62% for one-step and two-step designs, respectively. Similarly, the absence of the intention querying module results in a performance degradation of 11.25% and 21.25% for SiriusBI. This indicates that the dialogue analysis and intention querying modules are critical to the overall effectiveness of SiriusBI.
In addition, the ablation study reveals that SiriusBI one-step fine-tuned on a private dataset exhibits greater robustness compared to SiriusBI two-step. Concretely, the impact of removing intention querying on SiriusBI one-step is less pronounced, resulting in a performance drop of only 11.25%, compared to a more significant decline of 21.25% for SiriusBI two-step. The difference can be attributed to the fact that SiriusBI one-step acquires domain-specific knowledge from the training data during the learning process, which helps to mitigate the negative effects of lacking conversational understanding and intention querying.
@@ 5.3.2. Evaluation of Knowledge Management
Next, we evaluate the module using the dataset for knowledge retrieval with 7704 samples. In this evaluation, we compare SiriusBI against widely used baseline models BM25 (Robertson and Sparck Jones, 1976) and Vector (Karpukhin et al., 2020), both of which are prominent in the field of text retrieval. We assess the performance of these methods using the recall rates of knowledge retrieval for all keywords (KR) and schema linking (SL) for table columns only. Our results indicate that SiriusBI outperforms the baselines across both metrics, thereby ensuring that other modules leveraging knowledge management within SiriusBI can access accurate and relevant additional information.
| Baseline | KR Recall | SL Recall |
| BM25 | 90.28% | 89.58% |
| Vector | 93.06% | 93.06% |
| SiriusBI | 95.14% | 93.75% |
@@ 5.3.3. Data Insight Evaluations
Finally, we conduct experiments on the Data Insight module. Remind that the data insight dataset contains quantifiable questions involving comprehensive analysis and complex task planning.
| Baseline | Accuracy |
| Command-R+ | 75% |
| Qwen2.5 | 80% |
| GPT-4o | 83% |
| SiriusBI | 95% |
@@
For Comprehensive Analysis tasks, we use 100 real-world queries from the financial domain and conducts comparative experiments with three powerful LLM baselines: GPT-4o (Ope, [n.d.]), Qwen2.5 (Team, 2024) and Commnd R+ (cohere.ai Team, [n.d.]). The accuracy of responses to these queries serves as the evaluation criterion, and to ensure the reliability of the experimental results, each query is tested for three times. The specific experimental results are shown in Table 7. The findings indicate that the Data Insight module of SiriusBI maintains a notable performance advantage even against the currently most effective LLMs. The function call architecture (Kim et al., 2024) in SiriusBI enables it to outperform GPT-4o by 12% in Comprehensive Analysis tasks.
@@
For Complex Task Planning, we compare the Data Insight module with a method that only performs tool invocation on 50 high-quality queries. Compared to the simple tool invocation method, the multi-agent architecture adopted by SiriusBI shows a significant advantage, by improving the accuracy from 70% to 100% on this dataset. To facilitate understanding of the execution process of Complex Task Planning, we provide an example as follows:
6. Related Work
The integration of Artificial Intelligence (AI) and Machine Learning (ML) has fundamentally transformed data-driven decision-making and strategic planning within the realm of Business Intelligence (BI) (Rane et al., 2024). Machine Learning and Predictive Analytics are essential for enhancing BI, providing advanced tools for forecasting and identifying patterns in extensive datasets (Lee et al., 2022; Tamang et al., 2021). Additionally, Natural Language Processing (NLP) and Text Analytics enable the extraction of valuable insights from unstructured data, empowering businesses to understand customer sentiment, automate customer service, and enhance decision-making processes (Kaur et al., 2019; Ittoo et al., 2016). By leveraging the Internet of Things (IoT) and Big Data, businesses can increase their resilience and adaptability in an ever-evolving market (Alghamdi and Al-Baity, 2022; Klisarova-Belcheva et al., 2017). Advanced AI models, such as ChatGPT (Ope, [n.d.]), play a significantly augment BI by providing intelligent data interpretations and improving human-machine interactions, making data more accessible and comprehensible for non-technical users (Linkon et al., 2024).
A BI system typically stores vast amounts of information, such as customer and product sales data. In traditional BI systems, users need to use SQL to aggregate the required data and perform downstream analysis. Therefore, to lower the barrier to using BI capabilities and to create an intelligent BI system, NL2SQL has become a core task. Recently, Large Language Models (LLMs), including ChatGPT (Ope, [n.d.]), Codex (cod, [n.d.]) and Claude 3 (cla, [n.d.]), have demonstrated remarkable potential in addressing NL2SQL tasks. These advancements benefit from the introduction of prompt engineering (Giray, 2023), in-context learning (Dong et al., 2022), and the chain-of-thought approach (Wei et al., 2022). Research into prompt-based NL2SQL methods has not only provided new technical paradigms for addressing the challenges of this task but has also revealed opportunities for reducing technical costs in industrial applications. For instance, methods such as DIN-SQL (Pourreza and Rafiei, 2023), C3 (Dong et al., 2023), MAC-SQL (Wang et al., 2023), and SQL-PaLM (Sun et al., 2023), designed for models like ChatGPT and PaLM (Narang and Chowdhery, 2022), have demonstrated significant improvements in the performance on NL2SQL tasks.
Prior to the emergence of LLMs, solutions to the NL2SQL problem typically involved training an ”encoder-decoder” model in a sequence-to-sequence format (Cai et al., 2021; Cao et al., 2021; Choi et al., 2021; Gan et al., 2021; Wang et al., 2020), which heavily relied on a substantial amount of database and SQL-related data. The rise of large models has introduced new technical paradigms for traditional train-based methods. For example, Codes (Li et al., 2024) propose a training method tailored for the NL2SQL task, which enhances the performance of LLMs in this area. However, to achieve better performance in the NL2SQL task, some studies (Maamari et al., 2024; Pourreza et al., 2024; Caferoğlu and Ulusoy, 2024) opt to train LLMs such as ChatGPT (Ope, [n.d.]) and Gemini (gem, [n.d.]), which often comes with a high cost in practical applications.
In summary, although there have been pioneering efforts in academia to integrate LLMs into BI systems, these works have not fully addressed practical needs. From a practical perspective, we must also consider the flexible user interaction requirements, efficient NL2SQL capabilities, and a complete BI workflow. Therefore, how to incorporate LLMs into BI and develop a new generation of intelligent BI tools remains an open question.
@@ 7. Conclusion
In this paper, we propose SiriusBI, an deployed end-to-end LLM-based Business Intelligence system specifically designed to address three key challenges encountered in industrial scenarios. Our approach includes: 1) implementing multi-round dialogue with querying capabilities to overcome interaction limitations; 2) employing two SQL generation strategies to alleviate performance bottlenecks; and 3) establishing a comprehensive end-to-end workflow of the entire BI process to address functionality deficiencies. Currently, SiriusBI is deployed across Tencent’s finance, advertising and cloud sectors, serving dozens of enterprise clients. Extensive experiments on industrial datasets demonstrate that SiriusBI clearly outperforms competitive LLM-based baselines in both SQL generation and data insights generation.
@@ Acknowledgements.
We thank the researchers and engineers from the Data Platform department, Technology and Engineering Group (TEG) of Tencent for their technical and engineering support.References
@@
- Apa ([n.d.]) [n.d.]. Apache Superset. https://superset.apache.org/
- cla ([n.d.]) [n.d.]. Claude3. https://anthropic.com/claude
- gem ([n.d.]) [n.d.]. Gemini. https://gemini.google.com/
- Ope ([n.d.]) [n.d.]. Introduction of OpenAI Text Generation APIs. https://platform.openai.com/docs/guides/text-generation
- cod ([n.d.]) [n.d.]. OpenAI CodeX. https://openai.com/index/openai-codex
- Alghamdi and Al-Baity (2022) Noorah A. Alghamdi and Heyam H. Al-Baity. 2022. Augmented Analytics Driven by AI: A Digital Transformation beyond Business Intelligence. Sensors 22, 20 (2022), 8071.
@@
- Chen et al. (2024a) Xiaoxin Chen, Meng Wu, and Mangning Wang. 2024a. Application of business intelligence under deep neural network in credit scoring of bank users. Journal of Computational Methods in Sciences and Engineering 24, 3 (2024), 1585–1604.
@@
- Choudhury et al. (2023) Muntabir Hasan Choudhury, Lamia Salsabil, Himarsha R Jayanetti, Jian Wu, William A Ingram, and Edward A Fox. 2023. MetaEnhance: Metadata Quality Improvement for Electronic Theses and Dissertations of University Libraries. In 2023 ACM/IEEE Joint Conference on Digital Libraries. 61–65.
- cohere.ai Team ([n.d.]) cohere.ai Team. [n.d.]. The Cohere Platform. https://docs.cohere.com
@@
- Karpukhin et al. (2020) Vladimir Karpukhin, Barlas Oguz, Sewon Min, Patrick S. H. Lewis, Ledell Wu, Sergey Edunov, Danqi Chen, and Wen-tau Yih. 2020. Dense Passage Retrieval for Open-Domain Question Answering. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing. 6769–6781.
@@
- Klisarova-Belcheva et al. (2017) S Klisarova-Belcheva, G Ilieva, and T Yankova. 2017. BUSINESS INTELLIGENCE AND ANALYTICS–CONTEMPORARY SYSTEM MODEL. Trakia Journal of Sciences 15, 1 (2017), 298–304.
@@
- Li et al. (2024) Haoyang Li, Jing Zhang, Hanbing Liu, Ju Fan, Xiaokang Zhang, Jun Zhu, Renjie Wei, Hongyan Pan, Cuiping Li, and Hong Chen. 2024. CodeS: Towards Building Open-source Language Models for Text-to-SQL. Proceedings of the ACM on Management of Data 2, 3 (2024), 1–28.
@@
- Nuseir (2021) Mohammed T Nuseir. 2021. Designing business intelligence (BI) for production, distribution and customer services: a case study of a UAE-based organization. Business Process Management Journal 27, 4 (2021), 1275–1295.
@@
- Robertson and Sparck Jones (1976) SE Robertson and K Sparck Jones. 1976. Relevance Weighting of Search Terms. Journal of the American Society for Information Science 27, 3 (1976), 129–46.
@@
- Team (2024) Qwen Team. 2024. Qwen2.5: A Party of Foundation Models. https://qwenlm.github.io/blog/qwen2.5/
- team ([n.d.]) StarRocks team. [n.d.]. StarRocks. https://github.com/StarRocks/starrocks
- Wang et al. (2023) Bing Wang, Changyu Ren, Jian Yang, Xinnian Liang, Jiaqi Bai, Qian-Wen Zhang, Zhao Yan, and Zhoujun Li. 2023. MAC-SQL: A Multi-Agent Collaborative Framework for Text-to-SQL. arXiv preprint arXiv:2312.11242 (2023).
- Wang et al. (2020) Bailin Wang, Richard Shin, Xiaodong Liu, Oleksandr Polozov, and Matthew Richardson. 2020. RAT-SQL: Relation-Aware Schema Encoding and Linking for Text-to-SQL Parsers. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics. 7567–7578.
@@
- Yu et al. (2018) Tao Yu, Rui Zhang, Kai Yang, Michihiro Yasunaga, Dongxu Wang, Zifan Li, James Ma, Irene Li, Qingning Yao, Shanelle Roman, Zilin Zhang, and Dragomir R. Radev. 2018. Spider: A Large-Scale Human-Labeled Dataset for Complex and Cross-Domain Semantic Parsing and Text-to-SQL Task. In Proceedings of the 2018 Conference on Empirical Methods in Natural Language Processing. 3911–3921.

### v2

@@ 4.3.1. Evaluation of MRD-Q
In this part, we present the ablation study on dialogue analysis and intention querying using the MRD-Industry dataset. The experimental results are displayed in Table 3. Based on the experimental results, we can observe that, both the SiriusBI one-step and SiriusBI two-step heavily rely on the dialogue analysis and the intention querying modules. Specifically, when both modules are removed (without MRD-Q), the performance of SiriusBI declines significantly, with drops of 15.62% and 34.38% for one-step and two-step designs, respectively. Similarly, the absence of the intention querying module results in a performance degradation of 6.25% and 21.25% for SiriusBI. This indicates that the dialogue analysis and intention querying modules are critical to the overall effectiveness of SiriusBI.
@@
In addition, the ablation study reveals that SiriusBI one-step fine-tuned on a private dataset exhibits greater robustness compared to SiriusBI two-step. Concretely, the impact of removing intention querying in SiriusBI one-step is less pronounced, resulting in a performance drop of only 6.25%, compared to a more significant decline of 21.25% for SiriusBI two-step. The difference can be attributed to the fact that SiriusBI one-step acquires domain-specific knowledge from the training data during the learning process, which helps to mitigate the negative effects of lacking conversational understanding and intention querying.
4.3.2. Evaluation of Table Selection
While table selection directly impacts the UEX of SQL selection, we conduct an ablation study on this component. We use the queries in the SRD dataset and evaluate the performance using the Recall@5 metric as described before. The experimental results are shown in Table 4.
Baseline | Recall@5 | |
SiriusBI w/o Embed & Heat | 77.47% | -13.74% |
SiriusBI w/o Heat | 79.12% | -12.01% |
SiriusBI w/o Embed | 83.52% | -7.69% |
SiriusBI | 91.21% | - |
It is clear that both the embedding and heat information enhance the recall performance of table selection, which indicates the effectiveness of the combined ranking function in Equation 2.
@@ 4.3.3. Evaluation of Knowledge Management
Next, we evaluate the module using a dataset for knowledge retrieval with 7704 samples. In this evaluation, we compare SiriusBI against two widely used baseline models: BM25 (Robertson and Sparck Jones, 1976) and Vector (Karpukhin et al., 2020), both of which are prominent in the field of text retrieval. We assess the performance of these methods using two key metrics: KR Recall measuring the recall rate for retrieving relevant knowledge entries across all keywords and SL Recall measuring the recall rate specifically for linking table columns. The experimental results demonstrate the following performance: BM25 achieves a KR Recall of 90.28% and an SL Recall of 89.58%; Vector achieves a KR Recall of 93.06% and an SL Recall of 93.06%; SiriusBI achieves a KR Recall of 95.14% and an SL Recall of 93.75%. SiriusBI outperforms the baselines across both metrics, thereby ensuring that other modules leveraging knowledge management within SiriusBI can access accurate and relevant additional information.
@@ 4.3.4. Data Insight Evaluations
Finally, we conduct experiments on the Data Insight module. Remind that the data insight dataset contains 80 quantifiable questions involving comprehensive analysis and complex task planning.
@@
For evaluation, we use real-world queries from the financial domain and conduct comparative experiments with three powerful LLM baselines: GPT-4o (team, 2022), Qwen2.5 (Hui et al., 2024) and Commnd R+ (cohere.ai Team, 2024). The accuracy of responses to these queries serves as the evaluation criterion, and to ensure the reliability of the experimental results, each query is tested three times.
The results demonstrate that the Data Insight module of SiriusBI achieves an accuracy of 95%, significantly outperforming the baseline models. Specifically, Command-R+ achieves an accuracy of 75%, Qwen2.5 achieves 80%, and GPT-4o achieves 83%. These findings indicate that the Data Insight module of SiriusBI maintains a notable performance advantage even against the most effective LLMs currently in use. This superior performance can be attributed to the function call architecture in SiriusBI, which enables it to outperform GPT-4o by 12% in comprehensive data analysis tasks.
@@
Specifically, for complex task planning, we compare the Data Insight module with a method that only performs tool invocation on the aforementioned high-quality queries. Compared to the simple tool invocation method, the MRD-agent architecture adopted by SiriusBI shows a significant advantage, improving the accuracy from 70% to 100% on this dataset.
4.4. Case Study
In this part, we use an exemplar to demonstrate the whole workflow through a BI interaction. At the beginning, the user raises a question – What is the income of the Company A in 2024? The Dialogue Analysis module evaluates the semantic integrity of this query. In fact, this query contains both a “dimension” (2024) and a “metric” (income), making it semantically complete. The query then proceeds to the second stage.
The Intention Querying module recalls relevant information from the knowledge base based on the input query. Specifically, a piece of retrieved knowledge is a triplet – (company, by default, without special instructions, refers to Tencent, “What is gross profit of the company in 2023?”), which provides an explanation for the term “company”. The Intention Querying module further prompts the user for clarification as there still exists ambiguity within the query. In this case, the word “income” is ambiguous (with tax or not), the user will be asked whether it refers to the column “shouldincome” or “shouldincome_after”.
Afterwards, the SQL generation module selects the appropriate tables and translates the user query into an SQL query. The generated SQL statement is SELECT SUM(shouldincome_after) AS total_income FROM revenue_by_quarter WHERE YEAR
(ftime) = 2024 AND cname = Company A.
The Data Insight module receives the SQL queries and their corresponding execution results. Based on the user’s additional insight requirements like “Perform dimension attribution analysis on the revenue of Company A in 2024”, the module analyzes the user’s intent, decomposes tasks through the Planner, performs data preparation and invokes attribution tools, and finally generates an insight report, as shown in Figure 8.
4.5. Utility and Usability Study
To evaluate whether SiriusBI effectively improves users’ productivity and experience in real-world scenarios, we conducted a gray-release test within a business analysis team at Tencent data platform. This study aimed to quantify the system’s impact on operational efficiency and gather subjective feedback on usability. We randomly divided 40 business analysts into two groups – Group A (20 analysts): granted access to SiriusBI for daily tasks; Group B (20 analysts): continued using legacy tools. The experiment lasted for two weeks, during which we monitored both groups’ workflows. To ensure comparability, we controlled that tasks were distributed evenly across groups based on complexity and historical completion times. We find that Group A exhibited a 30.2% reduction in average task completion time compared to Group B. This efficiency gain is attributed to SiriusBI’s integrated workflow and advanced technical design. What’s more, Group A provided an average SUS score of 84.2/100 (SD = 6.5), categorizing SiriusBI as “excellent” according to the SUS benchmark (Brooke et al., 1996). Key strengths highlighted in feedback included: (1) 89% of users reported that domain-specific knowledge grounding reduced manual schema lookup time. (2) 92% agreed that SiriusBI’s data insight module “provided reliable support for business decision-making” (e.g., identifying root causes of revenue fluctuations). The gray-release test confirms that SiriusBI enhances both productivity and user experience by addressing the functionality deficiencies.
@@ 5. Conclusion
In this paper, we propose SiriusBI, a LLM-powered solution for business intelligence, specifically designed to address three key challenges encountered in industrial business intelligence scenarios. Our system 1) establishes a comprehensive end-to-end workflow of the entire business intelligence process to address functionality deficiencies; 2) implements multi-round dialogue with querying capabilities to overcome interaction limitations; 3) adopts a data-conditioned SQL generation method selection strategy to alleviate the great cost brought by cross-domain migration in SQL generation. Currently, SiriusBI is deployed across Tencent’s finance, advertising, and cloud sectors, serving enterprise clients with various profiles. Extensive experiments on public benchmark and industrial datasets demonstrate that SiriusBI clearly outperforms competitive LLM-based baselines in both SRD and MRD-based NL2SQL tasks. Detailed analysis on different modules demonstrate the rationality of our design. User studies further confirm that SiriusBI enhances both productivity and user experience.
@@ Acknowledgements.
We thank the engineers from the Data Platform department, Technology and Engineering Group (TEG) of Tencent for their technical and engineering support.References
@@
- Affolter et al. (2019) Katrin Affolter, Kurt Stockinger, and Abraham Bernstein. 2019. A comparative survey of recent natural language interfaces for databases. The VLDB Journal (2019).
- Bhagwan et al. (2014) Ranjita Bhagwan, Rahul Kumar, Ramachandran Ramjee, George Varghese, Surjyakanta Mohapatra, Hemanth Manoharan, and Piyush Shah. 2014. Adtributor: Revenue debugging in advertising systems. In 11th USENIX Symposium on Networked Systems Design and Implementation (NSDI 14).
- Box et al. (2015) George EP Box, Gwilym M Jenkins, Gregory C Reinsel, and Greta M Ljung. 2015. Time series analysis: forecasting and control. John Wiley & Sons.
- Brooke et al. (1996) John Brooke et al. 1996. SUS-A quick and dirty usability scale. Usability evaluation in industry (1996).
@@
- Chaudhuri et al. (2011) Surajit Chaudhuri, Umeshwar Dayal, and Vivek Narasayya. 2011. An overview of business intelligence technology. Commun. ACM (2011).
@@
- Chen et al. (2024c) Jianlv Chen, Shitao Xiao, Peitian Zhang, Kun Luo, Defu Lian, and Zheng Liu. 2024c. Bge m3-embedding: Multi-lingual, multi-functionality, multi-granularity text embeddings through self-knowledge distillation. arXiv preprint arXiv:2402.03216 (2024).
- Chen et al. (2024a) Xiaoxin Chen, Meng Wu, and Mangning Wang. 2024a. Application of business intelligence under deep neural network in credit scoring of bank users. Journal of Computational Methods in Sciences and Engineering (2024).
- Cheng et al. (2009) Hilary Cheng, Yi-Chuan Lu, and Calvin Sheu. 2009. An ontology-based business intelligence application in a financial knowledge management system. Expert Systems with Applications (2009).
@@
- cohere.ai Team (2024) cohere.ai Team. 2024. The Cohere Platform. https://docs.cohere.com
- Devlin et al. (2019) Jacob Devlin, Ming-Wei Chang, Kenton Lee, and Kristina Toutanova. 2019. Bert: Pre-training of deep bidirectional transformers for language understanding. In Proceedings of the 2019 conference of the North American chapter of the association for computational linguistics: human language technologies, volume 1 (long and short papers).
@@
- Floratou et al. (2024) Avrilia Floratou, Fotis Psallidas, Fuheng Zhao, Shaleen Deep, Gunther Hagleither, Wangda Tan, Joyce Cahoon, Rana Alotaibi, Jordan Henkel, Abhik Singla, et al. 2024. NL2SQL is a solved problem… Not!. In CIDR.
@@
- Gao et al. (2024) Yingqi Gao, Yifu Liu, Xiaoxia Li, Xiaorong Shi, Yin Zhu, Yiming Wang, Shiqi Li, Wei Li, Yuntao Hong, Zhiling Luo, Jinyang Gao, Liyu Mou, and Yu Li. 2024. XiYan-SQL: A Multi-Generator Ensemble Framework for Text-to-SQL. arXiv preprint arXiv:2411.08599 (2024).
@@
- Hui et al. (2024) Binyuan Hui, Jian Yang, Zeyu Cui, Jiaxi Yang, Dayiheng Liu, Lei Zhang, Tianyu Liu, Jiajun Zhang, Bowen Yu, Kai Dang, et al. 2024. Qwen2. 5-Coder Technical Report. arXiv preprint arXiv:2409.12186 (2024).
- Hundman et al. (2018) Kyle Hundman, Valentino Constantinou, Christopher Laporte, Ian Colwell, and Tom Soderstrom. 2018. Detecting spacecraft anomalies using lstms and nonparametric dynamic thresholding. In Proceedings of the 24th ACM SIGKDD international conference on knowledge discovery & data mining.
@@
- Karpukhin et al. (2020) Vladimir Karpukhin, Barlas Oguz, Sewon Min, Patrick S. H. Lewis, Ledell Wu, Sergey Edunov, Danqi Chen, and Wen-tau Yih. 2020. Dense Passage Retrieval for Open-Domain Question Answering. In Proceedings of the 2020 Conference on Empirical Methods in Natural Language Processing.
@@
- Ke et al. (2024) Wenjun Ke, Ziyu Shang, Zhizhao Luo, Peng Wang, Yikai Guo, Qi Liu, and Yuxuan Chen. 2024. Unveiling factuality and injecting knowledge for LLMs via reinforcement learning and data proportion. Science China Information Sciences (2024).
@@
- Kurt Schlegel (2023) David Pidsley Julian Sun Georgia O’Callaghan Christopher Long Kevin Quinn Fay Fei Edgar Macari Jamie O’Brien Kurt Schlegel, Anirudh Ganeshan. 2023. Magic Quadrant for Analytics and Business Intelligence Platforms. https://www.gartner.com/doc/reprints?id=1-2HW1JC8Q&ct=240620
@@
[…]
- Li et al. (2024b) Chaofan Li, Yingxia Shao, and Zheng Liu. 2024b. SEA-SQL: Semantic-Enhanced Text-to-SQL with Adaptive Refinement. arXiv preprint arXiv:2408.04919 (2024).
@@
- Li et al. (2024c) Haoyang Li, Jing Zhang, Hanbing Liu, Ju Fan, Xiaokang Zhang, Jun Zhu, Renjie Wei, Hongyan Pan, Cuiping Li, and Hong Chen. 2024c. CodeS: Towards Building Open-source Language Models for Text-to-SQL. Proceedings of the ACM on Management of Data 2, 3 (2024), 1–28.
@@
- Li (2024) Jian-Zhong Li. 2024. opengauss: An open-source database for the era of artificial intelligence. Journal of Computer Science and Technology (2024).
- Li et al. (2019) Zeyan Li, Chengyang Luo, Yiwei Zhao, Yongqian Sun, Kaixin Sui, Xiping Wang, Dapeng Liu, Xing Jin, Qi Wang, and Dan Pei. 2019. Generic and robust localization of multi-dimensional root causes. In 2019 IEEE 30th International Symposium on Software Reliability Engineering (ISSRE).
@@
- Liu et al. (2024) Xinyu Liu, Shuyu Shen, Boyan Li, Peixian Ma, Runzhi Jiang, Yuyu Luo, Yuxin Zhang, Ju Fan, Guoliang Li, and Nan Tang. 2024. A Survey of NL2SQL with Large Language Models: Where are we, and where are we going? arXiv preprint arXiv:2408.05109 (2024).
- Lundberg and Lee (2017) Scott M Lundberg and Su-In Lee. 2017. A unified approach to interpreting model predictions. Advances in neural information processing systems 30 (2017).
@@
- Meduri et al. (2021) Venkata Vamsikrishna Meduri, Abdul Quamar, Chuan Lei, Vasilis Efthymiou, and Fatma Ozcan. 2021. BI-REC: guided data analysis for conversational business intelligence. arXiv preprint arXiv:2105.00467 (2021).
@@
- Nuseir (2021) Mohammed T Nuseir. 2021. Designing business intelligence (BI) for production, distribution and customer services: a case study of a UAE-based organization. Business Process Management Journal (2021).
@@
- Radford et al. (2019) Alec Radford, Jeffrey Wu, Rewon Child, David Luan, Dario Amodei, Ilya Sutskever, et al. 2019. Language models are unsupervised multitask learners. OpenAI blog (2019).
@@
- Reimers and Gurevych (2019) Nils Reimers and Iryna Gurevych. 2019. Sentence-BERT: Sentence Embeddings using Siamese BERT-Networks. In Proceedings of the 2019 Conference on Empirical Methods in Natural Language Processing and the 9th International Joint Conference on Natural Language Processing (EMNLP-IJCNLP).
@@
- Robertson and Sparck Jones (1976) SE Robertson and K Sparck Jones. 1976. Relevance Weighting of Search Terms. Journal of the American Society for Information Science (1976).
@@
- Sun et al. (2019) Yu Sun, Shuohuan Wang, Yukun Li, Shikun Feng, Xuyi Chen, Han Zhang, Xin Tian, Danxiang Zhu, Hao Tian, and Hua Wu. 2019. Ernie: Enhanced representation through knowledge integration. arXiv preprint arXiv:1904.09223 (2019).
@@
- Team (2024) Claude Team. 2024. Claude3. https://anthropic.com/claude
- Team et al. (2023) Gemini Team, Rohan Anil, Sebastian Borgeaud, Jean-Baptiste Alayrac, Jiahui Yu, Radu Soricut, Johan Schalkwyk, Andrew M Dai, Anja Hauth, Katie Millican, et al. 2023. Gemini: a family of highly capable multimodal models. arXiv preprint arXiv:2312.11805 (2023).
- team (2022) OpenAI team. 2022. Introduction of OpenAI Text Generation APIs. https://platform.openai.com/docs/guides/text-generation
- team (2023a) OpenAI team. 2023a. OpenAI CodeX. https://openai.com/index/openai-codex
- team (2023b) StarRocks team. 2023b. StarRocks. https://github.com/StarRocks/starrocks
- Wang et al. (2023b) Bing Wang, Changyu Ren, Jian Yang, Xinnian Liang, Jiaqi Bai, Qian-Wen Zhang, Zhao Yan, and Zhoujun Li. 2023b. MAC-SQL: A Multi-Agent Collaborative Framework for Text-to-SQL. arXiv preprint arXiv:2312.11242 (2023).
- Wang et al. (2020a) Bailin Wang, Richard Shin, Xiaodong Liu, Oleksandr Polozov, and Matthew Richardson. 2020a. RAT-SQL: Relation-Aware Schema Encoding and Linking for Text-to-SQL Parsers. In Proceedings of the 58th Annual Meeting of the Association for Computational Linguistics. 7567–7578.
@@
- Wang et al. (2025) Xin Wang, Zirui Chen, Haofen Wang, Leong Hou U, Zhao Li, and Wenbin Guo. 2025. Large language model enhanced knowledge representation learning: A survey. Data Science and Engineering (2025).
- Wang et al. (2023a) Yizhong Wang, Yeganeh Kordi, Swaroop Mishra, Alisa Liu, Noah A Smith, Daniel Khashabi, and Hannaneh Hajishirzi. 2023a. Self-Instruct: Aligning Language Models with Self-Generated Instructions. In Proceedings of the 61st Annual Meeting of the Association for Computational Linguistics (Volume 1: Long Papers).
- Wang et al. (2020b) Yaqing Wang, Quanming Yao, James T Kwok, and Lionel M Ni. 2020b. Generalizing from a few examples: A survey on few-shot learning. ACM computing surveys (csur) (2020).
@@
- Weir et al. (2020) Nathaniel Weir, Prasetya Utama, Alex Galakatos, Andrew Crotty, Amir Ilkhechi, Shekar Ramaswamy, Rohin Bhushan, Nadja Geisler, Benjamin Hättasch, Steffen Eger, et al. 2020. Dbpal: A fully pluggable nl2sql training pipeline. In Proceedings of the 2020 ACM SIGMOD International Conference on Management of Data.
@@
- Wu et al. (2021) Kun Wu, Lijie Wang, Zhenghua Li, Ao Zhang, Xinyan Xiao, Hua Wu, Min Zhang, and Haifeng Wang. 2021. Data augmentation with hierarchical SQL-to-question generation for cross-domain text-to-SQL parsing. arXiv preprint arXiv:2103.02227 (2021).
- Yang et al. (2024) An Yang, Baosong Yang, Beichen Zhang, Binyuan Hui, Bo Zheng, Bowen Yu, Chengyuan Li, Dayiheng Liu, Fei Huang, Haoran Wei, et al. 2024. Qwen2. 5 technical report. arXiv preprint arXiv:2412.15115 (2024).
- Yao et al. (2023) Shunyu Yao, Jeffrey Zhao, Dian Yu, Nan Du, Izhak Shafran, Karthik Narasimhan, and Yuan Cao. 2023. React: Synergizing reasoning and acting in language models. In International Conference on Learning Representations (ICLR).
@@
- Zhang et al. (2023b) Baoli Zhang, Haining Xie, Pengfan Du, Junhao Chen, Pengfei Cao, Yubo Chen, Shengping Liu, Kang Liu, and Jun Zhao. 2023b. ZhuJiu: A Multi-dimensional, Multi-faceted Chinese Benchmark for Large Language Models. In Proceedings of the 2023 Conference on Empirical Methods in Natural Language Processing: System Demonstrations.
- Zhang et al. (2024) Jiayu Zhang, Jiangrong Shen, Zeke Wang, Qinghai Guo, Rui Yan, Gang Pan, and Huajin Tang. 2024. SpikingMiniLM: energy-efficient spiking transformer for natural language understanding. Science China Information Sciences (2024).
- Zhang et al. (2023a) Yi Zhang, Jan Deriu, George Katsogiannis-Meimarakis, Catherine Kosten, Georgia Koutrika, and Kurt Stockinger. 2023a. Sciencebenchmark: A complex real-world benchmark for evaluating natural language to sql systems. arXiv preprint arXiv:2306.04743 (2023).
- Zhou et al. (2024) Xuanhe Zhou, Zhaoyan Sun, and Guoliang Li. 2024. Db-gpt: Large language model meets database. Data Science and Engineering (2024).